OpenStudy (anonymous):

Given f(x) = 3x + 2 and g(x) = x^2 + 5x - 2, find f(x) - g(x). Please help

OpenStudy (anonymous):

f(x)-g(x)=3x+2-(x^2+5x-2) =3x+2-x^2-5x+2 =-2x-x^2+4 =-(x^2+2x-4)

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (anonymous):

welcome =)